Account Current Year Prior Year
Net sales (all credit) $445,400 $362,000
Costs of goods sold $220,000 $185,000
Gross profit $225,400 $177,000
Income from operations $72,000 $80,000
Interest expense $8,000 $14,000
Net income $40,000 $35,000
Cash $34,000 $28,000
Accounts receivable, net $42,000 $62,000
Inventory $120,000 $100,000
Prepaid expenses $4,000 $2,000
Total current assets $200,000 $192,000
Total long-term assets $200,000 $230,000
Total current liabilities $100,000 $165,000
Total long-term liabilities $44,000 $95,000
Common stock, no par, $8,000 shares, value $120/share $80,000 $80,000

What balance sheet account changed the most from last year to this year? (as a %) Show work + explanation

What income statement account changed the most from last year to this year? (as a %) Show work + explanation
Answer #1

Change in Current Year (A) Prior Year (B) Amount (C) = (A)-(B) Percentage (D) = (C)/(B) $ 445,400 S 362.000 $ 83,400 23.04% $